OpenStudy (anonymous):

Find the derivative of cos(arcsin 2x)

OpenStudy (luigi0210):

You're gonna have to use the chain rule on this: \[\large \frac{d}{dx}cos(arcsin2x)=(cos(arcsin2x))'*(arcsin2x)'*(2x)'\]

OpenStudy (anonymous):

thank you!

OpenStudy (luigi0210):

Got it from here? Good luck!

OpenStudy (anonymous):

No haha i was saying thank you in advance haha

OpenStudy (anonymous):

\[\frac{ d }{ dx }\left( \cos x \right)=-\sin x\] \[\frac{ d }{ dx }\left( \sin^{-1} x \right)=\frac{ 1 }{\sqrt{1-x^2} }\]